# Quantifying sunspot group nesting with density-based unsupervised clustering

### v1

@@ 4 Discussion
First, the typical scale and temporal persistence of nests identified here align well with the classical findings of cas86 and bro90. Nests commonly persist for one to several solar rotations and can extend across tens of degrees in Carrington longitude, consistent with the recurrent emergence of magnetic flux from long-lived subsurface concentrations. The observed displacement of nest centres in longitude, often independent of the local differential rotation, matches earlier descriptions of active-region complexes and reflects underlying magnetic structures that rotate with their own characteristic phase speeds.
@@
An additional result comes from analysing the RGO and KMAS catalogues independently over their overlap period. RGO consistently yields slightly higher nesting degrees, which we attribute to its more complete sampling of small sunspot groups. Despite this offset, both datasets display the same latitude dependence, similar ranges of $D$, and comparable cycle-to-cycle variability. This agreement indicates that the clustering patterns recovered by our method represent intrinsic properties of solar emergence rather than artefacts of a particular catalogue.
@@
Third, the synthetic-data validation demonstrates that the chosen clustering parameters recover the true nesting degree with $\lesssim 10$% scatter. Importantly, the behaviour of DBSCAN differs from classical active-longitude analyses, which typically rely on sinusoidal fits or phase-folding of emergence patterns. DBSCAN captures clustering of arbitrary shape, enabling us to detect both compact nests and extended, drifting structures that would be difficult to characterise using traditional methods.
@@
A further result is the moderate but clear positive correlation between nesting degree and overall cycle activity (Section 3.4). Stronger cycles, characterised by larger emergent flux, tend to show higher clustering. This behaviour is consistent with the idea that more coherent toroidal flux systems during strong cycles feed repeated localised flux emergence, whereas weaker cycles exhibit reduced magnetic organisation. It is also consistent with a similar possibility proposed for solar-type stars, where an increase in the nesting degree with the activity level could explain the observed trends in photometric variability amplitudes (isik20).
@@
Finally, these results have implications for Sun-as-a-star variability studies. If more than half of all sunspot groups emerge within nests, as our analysis indicates, then the rotational modulation of solar irradiance is strongly influenced by nest persistence and longitudinal asymmetry. This connects directly to stellar observations, where active longitudes and long-lived spot complexes are commonly inferred in photometric time series.
@@
The quantitative nesting degree ($D\approx 0.61$) provides a benchmark for comparison with stellar active-region nesting (isik20). If similar nesting fractions occur on other Sun-like stars, photometric variability and rotational modulation amplitudes should reflect both the instantaneous distribution of active regions and their tendency to cluster in spatial-temporal complexes.
A number of limitations should be noted. DBSCAN relies on a single spatial–temporal scale parameter, which cannot capture variations in nest size within an analysis window. Our approach also treats longitude and time symmetrically after standardisation, whereas the physical processes governing flux emergence may introduce anisotropy. Furthermore, the longitude–time projection does not account for potential clustering in latitude. Future work could incorporate full 3D clustering or probabilistic models that allow continuous tracking of nest evolution.
@@ 5 Conclusions
  • Nesting degree correlates positively with overall cycle strength, suggesting that more coherent toroidal flux bundles sustain repeated flux emergence.
@@
  • Despite quantitative differences, RGO and KMAS yield qualitatively consistent nesting picture, suggesting the robustness of our approach and the results. Quantitative differences are at least partially due to non-exhaustive observational coverage.

### v2

@@ 4 Discussion
First, the typical scale and temporal persistence of nests identified here align well with the classical findings of Castenmiller et al. (1986) and Brouwer and Zwaan (1990). Nests commonly persist for one to several solar rotations and can extend across tens of degrees in Carrington longitude, consistent with the recurrent emergence of magnetic flux from long-lived subsurface concentrations. The observed displacement of nest centres in longitude, often independent of the local differential rotation, matches earlier descriptions of active-region complexes and reflects underlying magnetic structures that rotate with their own characteristic phase speeds.
@@
Recent studies using density-based clustering identified persistent active longitudes ($\sim 40^{\circ}$ wide, 15-25 Carrington rotations) by analysing all latitudes together (Korsós et al., 2024; Császár et al., 2026). Our latitude-sliced approach quantifies nesting in both longitude and latitude over full solar cycles, revealing strong mid-latitude enhancement but no persistent longitudinal preference (Fig. 7). This cycle-scale longitudinal homogeneity reflects differential rotation dispersing shorter-lived active longitude structures across all longitudes over decadal periods.
@@
The third result comes from analysing the RGO and KMAS catalogues independently over their overlap period. RGO consistently yields slightly higher nesting degrees, which we attribute to its more complete sampling of small sunspot groups. Despite this offset, both datasets display the same latitude dependence, similar ranges of $D$, and comparable cycle-to-cycle variability. This agreement indicates that the clustering patterns recovered by our method represent intrinsic properties of solar emergence rather than artefacts of a particular catalogue.
@@
Fourth, the synthetic-data validation demonstrates that the chosen clustering parameters recover the true nesting degree with $\lesssim 10$% scatter. Importantly, the behaviour of DBSCAN differs from classical active-longitude analyses, which typically rely on sinusoidal fits or phase-folding of emergence patterns. DBSCAN captures clustering of arbitrary shape, enabling us to detect both compact nests and extended, drifting structures that would be difficult to characterise using traditional methods.
@@
A further result is the moderate but clear positive correlation between nesting degree and overall cycle activity (Section 3.4). Stronger cycles, characterised by larger emergent flux, tend to show higher clustering. This behaviour is consistent with the idea that more coherent toroidal flux systems during strong cycles feed repeated localised flux emergence, whereas weaker cycles exhibit reduced magnetic organisation. It is also consistent with a similar possibility proposed for solar-type stars, where an increase in the nesting degree with the activity level could explain the observed trends in photometric variability amplitudes (Işık et al., 2020).
@@
Finally, these results have implications for Sun-as-a-star variability studies. If more than half of all sunspot groups emerge within nests, as our analysis indicates, then the rotational modulation of solar irradiance is strongly influenced by nest persistence and longitudinal asymmetry. This connects directly to stellar observations, where active longitudes and long-lived spot complexes are commonly inferred in photometric time series (Özavcı et al., 2018; Breton et al., 2024). The quantitative nesting degree ($D\approx 0.61$) provides a benchmark for comparison with stellar active-region nesting (Işık et al., 2020). If similar nesting fractions occur on other Sun-like stars, photometric variability and rotational modulation amplitudes should reflect both the instantaneous distribution of active regions and their tendency to cluster in spatial-temporal complexes.
A number of general limitations should be noted. DBSCAN relies on a single spatial–temporal scale parameter, which cannot capture variations in nest size within an analysis window. Our approach also treats longitude and time symmetrically after standardisation, whereas the physical processes governing flux emergence may introduce anisotropy. Furthermore, the longitude–time projection does not account for potential clustering in latitude. Future work could incorporate full 3D clustering or probabilistic models that allow continuous tracking of nest evolution.
The inter-nest length scale $L_{\rm typ}$ (Sect. 3.5) represents a phase-space density measure integrating over full cycle durations and uniform spatial distribution across $360^{\circ}\times 5^{\circ}$ bands. Consequently, $L_{\rm typ}$ does not reflect actual local spacing between neighbouring nests within active zones, where emergence concentrates preferentially (Gaizauskas et al., 1983). Nevertheless, the robust anti-correlation with activity ($r\approx-0.65$, $p<10^{-13}$) reveals a significant pattern: during solar maximum, nests pack into denser configurations, approaching typical SG elongations. This convergence may reflect hierarchical magnetic organisation, with compact nests clustering into larger composite structures spanning up to $55^{\circ}$ longitude (Brouwer and Zwaan, 1990). Refined analysis focusing on active emergence periods rather than cycle-integrated densities would better characterise local inter-nest spacing but lies beyond the scope of the present study.
@@ 5 Conclusions
  • Nesting degree correlates positively with overall cycle strength, suggesting that more coherent toroidal flux bundles sustain repeated flux emergence. Importantly, this correlation strengthens when small groups are progressively excluded, indicating that nesting represents genuine clustering in flux emergence rather than statistical artefacts from abundant small groups.
@@
  • Nesting exhibits hierarchical structure: small sunspot groups cluster strongly around large active regions (nesting degree $\sim$0.5–0.6), while large groups emerge with greater spatial independence (asymptotic nesting degree $\sim$0.25).
  • Characteristic inter-nest spacing shows strong anti-correlation with activity, contracting from $\sim$200–500 Mm during low activity to $\sim$60–100 Mm during solar maximum, comparable to typical sunspot group elongations (35–117 Mm) and indicating dense spatial packing of emergence sites.
  • Despite quantitative differences, RGO and KMAS yield qualitatively consistent nesting picture, suggesting the robustness of our approach. Quantitative differences are at least partially due to non-exhaustive observational coverage.
@@
{acks}
We thank the anonymous reviewer for their comments and suggestions.